OpenStudy (anonymous):

IT TAKES ROBERT 9 HRS.LONGRE TO CONSTRUCT A FENCE THAT ITR TAKES ELDRIN. IF THEY WORK TOGETHER THEY CAN CONSTRUCT THE FENCE IN 20 HRS. HOW LONG WOULD IT TAKE EACH WORKING ALONE TO CONSTRUCT A FENCE?

OpenStudy (anonymous):

Why are you yelling?

OpenStudy (anonymous):

I got 36 and 45 hours.

OpenStudy (anonymous):

Let eldrin complete the work in x hrs therefore robert completes it in x+9 hrs Work done by eldrin in 1 hr = 1/x Work done by robert in 1 hr= 1/x+9 Therefore work doen by elderin and robert together in a day = (1/x) +(1/x+9) By data work is done in 20 hrs, therefore [ (1/x) + (1/x+9)]*20=1 (total work done amounts to 1) Therefore simplifying we get = (20/x) + (20/x+9)=1 20(x+9) + 20x/x(x+9)=1 (taking LCM) 20x + 180 + 20x / x^2 +9x = 1 40x +180=x^2 +9x (cross multiplication) x^2 -31x -180 = 0 x^2 -36x +5x -180=0 x(x-36) +5 (x-36)=0 (x+5)(x-36)=0 Therefore x=36(Applicable) or x=-5(Not applicable) Therefore eldrin takes 36 hrs and robert 45 hrs !!! Please give a good answer!!!

OpenStudy (anonymous):

this is how i also got 36 and 45

OpenStudy (anonymous):

@alexray19: By solving \[ x (9 + x) - 20 (9 + 2 x) = 0 \] then then neglecting the negative solution.